HUBBLE SPACE TELESCOPE DAILY REPORT #834 PERIOD COVERED: 0800 EST Daily Status Report as of 061/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FOS/RD 4123 (FOS Spectral Flat Fields Monitor Revised) The Faint Object Spectrograph (Red Side only) was used to make observations to standard star WD0501+527 to make standard spectral time dependent ultraviolet flat field calibrations. All of the observations were received and displayed in the OSS. They were reported to look okay. 1.2 Completed PC 4712 (WF/PC Calibration: Measles Monitor) The Planetary Camera was used to make six Earth calibration observations using filter F517N, both with and without filter F122M. The goal is to monitor the appearance and the number of measles (i.e., a form of contamination). Of the six images which were received, the fifth was reported to be underexposed, and the others were okay. 1.3 Completed WFC/FOS/BL 3707 (HST Observations of Periodic Comets) The Wide Field Camera and the Faint Object Spectrograph (On both the Red and Blue sides) were used to study periodic comet Schaumasse. A decision observation was made, and a set of corrective offsets were uplinked to the spacecraft. No problems were reported. The spectral observations occurred during LOS, and still remain on the science tape recorder. They will be analyzed later. 1.4 Completed PC 4093 (Survey of Nuclear Regions of 102 Markarian Galaxies) The Planetary Camera was used to make images of the inner regions of the Markarian galaxy MARK96. The observations were made in course track. The observations were monitored in the OSS, and no problems were reported. 1.5 Completed PC 4027 (Non-Proprietary Snapshots Survey Round 2) The Planetary Camera was used to make a non-proprietary snapshot image of the quasar US1420. The observations was executed as scheduled, and no problems were reported. 1.6 Completed WFC 3172 (WF/PC SV Observations: Internal Monitor Proposal) The Wide Field Camera was used to make an internal bias calibration image. No external targets were used. The observations were executed as scheduled, and monitored in the OSS. No problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 8 Successful : 8 Scheduled Reacquisitions : 5 Successful : 5 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 13 Successful : 13 2.3 Significant HSTARs: None 2.4 Operations Note: None 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete WFC 4658 (Determination of Extragalactic distance Scale) 3.2 Complete Two Sets of PC 3603 (Snapshot Of Protoplanetary Nebulae) 3.3 Complete PC 3877 (Crane, Augmentation, Stellar Astrophysics) 3.4 Complete S/C 4816 (FOC/48 Reactivation (Camera Section Only)) 3.5 Complete WFC 3172 (WF/PC SV Observations: Internal Monitor Proposal) 3.6 Complete WFC/FOS/BL 3707 (HST Observations of Periodic Comets) 3.7 Uplink Updated Ephemeris Tables, per MOSES 2871 3.8 Initiate FOC/96 4071 (Ultraviolet Studies of Galaxies) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: The GSFC Simulation (Type 5 - Deployment) is scheduled for Thursday March 4, 1993 (063/1300Z-064/0100Z).